OpenStudy (anonymous):

Angie was working on solving the exponential equation 23^x = 6; however she isn't sure where to start. Describe to Angie how to solve this equation and how solving it would be different if the bases were equal. The only thing I need help with is how solving it would be different if the bases were equal

OpenStudy (unanimoose):

If the bases were equal you would just need to compare the exponent.

OpenStudy (anonymous):

Okay. How would that look?

OpenStudy (unanimoose):

Okay say you had: \[9^x = 16\]

OpenStudy (unanimoose):

Sorry the site lagged out for a sec.

OpenStudy (unanimoose):

Ok so we can rewrite this as: \[3^3x = 4^2\]

OpenStudy (unanimoose):

Woops, hold on.

OpenStudy (unanimoose):

Let me use a different example.

OpenStudy (unanimoose):

\[64^x = 16\]

OpenStudy (unanimoose):

Ok, we can rewrite this as: \[4^3x = 4^2\]

OpenStudy (unanimoose):

That x is meant to be part of the exponent but Latex didn't let me do that.

OpenStudy (unanimoose):

So since the bases are the same we can just compare the exponent and just solve for: 3x = 2 Since the bases are equivalent and nothing needs to be done with them. @Tarynrachelle

OpenStudy (unanimoose):

Hope this helps.

OpenStudy (anonymous):

Okay.. Let me see if I get it haha. Thank you :)

OpenStudy (unanimoose):

Np. :)